OpenStudy (anonymous):

WILL MEDAL!!!! A company sells cardboard that is nine and one-eighth millimeters thick. Write an equation to solve for the number of sheets of cardboard, c, in a pile that is two hundred fifty-five and one-half millimeters thick. What must be done to both sides to solve the equation?

OpenStudy (anonymous):

@HeroOfLove111 @Hero @amistre64 @highschoolmom2010 @boricua5 @ninaesb

OpenStudy (amistre64):

what equation do you come up with?

OpenStudy (anonymous):

i didnt i dont understand it

OpenStudy (amistre64):

spose we work with simpler numbers and a different medium. if you have 1 dollar that is worth $1, how many dollars would it take to have $500? how would you solve this?

OpenStudy (anonymous):

dont you multiply 1 by 500 and you would need to get 500 $1 bills in order to have 500$

OpenStudy (amistre64):

yes, in this case we have cardboard that is equal to 9 1/8 and we want to find how many cardboards equal up to get 255 1/2 so same concept, 9 1/8, times 'c', is equal to 255 1/2

OpenStudy (anonymous):

so it would be the 255 1/2???

OpenStudy (amistre64):

are the fraction bad for you?

OpenStudy (anonymous):

yea they are

OpenStudy (amistre64):

just think of them as a number, any number to get the feel for it. in hindsight a 5 dollar bill might have been better to use as an example :) lets say c = 9, how many cs does it take to make 225?

OpenStudy (anonymous):

28.33333333333333

OpenStudy (amistre64):

\[\underbrace{c+c+c+c+...+c}_{n~times}=225\] but c=9 sooo \[\underbrace{9+9+9+9+...+9}_{n~times}=225\] or written another way:\[9n=225\] dividing both sides by 9 gets us the value of n, agreed?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then this is how our equation sets up, but with the fraction involved :) and by the way 225/9 = 25, not 28.333333

OpenStudy (amistre64):

so,\[(9\frac18)n=225\frac12\] divide both sides by 9 1/8 \[n=\frac{225\frac12}{9 \frac18}\]

OpenStudy (anonymous):

okay